OpenStudy (anonymous):

Answer please i'll give a medal. Find the quotient. 2/5 divide (-3/10)

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Multiply by the reciprocal.

OpenStudy (anonymous):

2/5(-10/3)=20/15=4/3

OpenStudy (anonymous):

-4/3